# Emergence of a High-Dimensional Abstraction Phase in Language Transformers

### v1

@@ H.1 Tasks
H.2 Setup
We use the training and test data provided by Conneau et al. [2018]. We train a MLP classifier for each task and each layer of each LM, repeating the experiment with 5 different seeds.
@@ H.2 Setup
For all other hyperparameters, we used the default values set by the the Scikit-learn library [Pedregosa et al., 2011].
We repeated all experiments after shuffling the example labels (both at training and test time). This provides a baseline for each task, as well as functioning as a sanity check that a probing classifier is not so powerful as to simply memorize arbitrary patterns in the representations [Hewitt and Liang, 2019].
H.3 Additional Results
Figure H.1 reports the probing performance for OLMo and Mistral, and for surface form (top row) as well as syntactic and semantic tasks (bottom row). As for Llama, OPT, and Pythia, we observe that the first ID peak converges to viable syntactic/semantic abstraction of inputs, while discarding information about surface form.
Figure H.2 shows that, on the shuffled corpora ablation, the MLP probes perform at chance. We then confirm that semantic and syntactic information is contained in the model representations and not in the probes.
Figure H.1: Linguistic knowledge probing performance ±$\pm 2$ SDs across 5 random seeds is plotted with the ID across layers for OLMo and Mistral (left to right). (Top row) Surface form tasks Sentence Length and Word Content, where probe performance decreases through the ID peak. (Bottom row) Semantic and syntactic tasks Bigram Shift, Coordination Inversion and Odd Man Out, where probe performance for all tasks attains maximum within the ID peak.
Figure H.2: Linguistic probe validation accuracy for semantic, syntactic, and surface tasks (solid lines) and their shuffled versions (dashed lines) are shown across layers for all models. The probing performance on shuffled corpora is constant at chance for all tasks and models.
Appendix I Downstream Tasks
I.1 Tasks
@@
We use a random balanced subset ($N=30588$) of Kaggle’s jigsaw toxic comment classification challenge [Adams et al., 2017], where each data point consists of a natural language comment and its binary toxicity label.
@@
We use a dataset of IMDb movie reviews [Maas et al., 2011], where each data point consists of a natural language film review and a corresponding label $\in\{\text{positive},\text{negative}\}$. To train the linear probes, we take a sample of $N=25000$ corresponding to the train split on HuggingFace. We repeat the experiment with 5 distinct seeds.
I.2 Setup
To train the linear probes, we first divide the data at random into train (80%) and validation (20%) sets. We feed each training set through each model and gather the last token hidden representations at each layer. Then, using PyTorch [Paszke et al., 2019], we train one linear probe per layer with hyperparameters as follows,
@@ I.2 Setup
I.3 Additional Results
Similar to Llama, OPT, and Pythia, Figure I.1 shows that validation performance for downstream tasks for Mistral and OLMo converge in the ID peak.
Figure I.1: Validation performance on downstream transfer tasks on toxicity and sentiment classification for Mistral (left) and OLMo (right). All validation accuracy curves are plotted with $\pm$ 2 standard deviations over 5 random seeds. For all models, classification validation accuracy converges within the ID peak.

### v2

@@ I.1 Tasks
I.2 Setup
We use the training and test data provided by Conneau et al. (2018). We train a MLP classifier for each task and each layer of each LM, repeating the experiment with 5 different seeds.
@@ I.2 Setup
For all other hyperparameters, we used the default values set by the the Scikit-learn library (Pedregosa et al., 2011).
We repeated all experiments after shuffling the example labels (both at training and test time). This provides a baseline for each task, as well as functioning as a sanity check that a probing classifier is not so powerful as to simply memorize arbitrary patterns in the representations (Hewitt & Liang, 2019). Note that, since, as shown in Figure I.2, performance is essentially constant (and at chance) on shuffled labels, accuracy is in our case equivalent to selectivity, the measure recommended by Hewitt & Liang (2019). We thus focus on accuracy, as the more familiar measure.
I.3 Additional Results
Figure I.1 reports the probing performance for OLMo and Mistral, and for surface form (top row) as well as syntactic and semantic tasks (bottom row). As for Llama, OPT, and Pythia, we observe that the first ID peak converges to viable syntactic/semantic abstraction of inputs, while discarding information about surface form.
Figure I.2 shows that, on the shuffled corpora ablation, the MLP probes perform at chance. We then confirm that semantic and syntactic information is contained in the model representations and not in the probes.
Figure I.1: Linguistic knowledge probing performance ±$\pm 2$ SDs across 5 random seeds is plotted with the ID across layers for OLMo and Mistral (left to right). (Top row) Surface form tasks Sentence Length and Word Content, where probe performance decreases through the ID peak. (Bottom row) Semantic and syntactic tasks Bigram Shift, Coordination Inversion and Odd Man Out, where probe performance for all tasks attains maximum within the ID peak.
Figure I.2: Linguistic probe validation accuracy for semantic, syntactic, and surface tasks (solid lines) and their shuffled versions (dashed lines) are shown across layers for all models. The probing performance on shuffled corpora is constant at chance for all tasks and models.
Appendix J Downstream Tasks
J.1 Tasks
@@ I.3 Additional Results
We use a random balanced subset ($N=30588$) of Kaggle’s jigsaw toxic comment classification challenge (Adams et al., 2017), where each data point consists of a natural language comment and its binary toxicity label.
@@
We use a dataset of IMDb movie reviews (Maas et al., 2011), where each data point consists of a natural language film review and a corresponding label $\in\{\text{positive},\text{negative}\}$. To train the linear probes, we take a sample of $N=25000$ corresponding to the train split on HuggingFace. We repeat the experiment with 5 distinct seeds.
J.2 Setup
To train the linear probes, we first divide the data at random into train (80%) and validation (20%) sets. We feed each training set through each model and gather the last token hidden representations at each layer. Then, using PyTorch (Paszke et al., 2019), we train one linear probe per layer with hyperparameters as follows,
@@
J.3 Additional Results
Similar to Llama, OPT, and Pythia, Figure J.1 shows that validation performance for downstream tasks for Mistral and OLMo converge in the ID peak.
Figure J.1: Validation performance on downstream transfer tasks on toxicity and sentiment classification for Mistral (left) and OLMo (right). All validation accuracy curves are plotted with $\pm$ 2 standard deviations over 5 random seeds. For all models, classification validation accuracy converges within the ID peak.